OpenStudy (anonymous):

How do you do matrix multiplication? 2 4 times -1 3 1 1 5 2

jimthompson5910 (jim_thompson5910):

The first row of the left matrix is 2 4 The first column of the right matrix is -1 5

jimthompson5910 (jim_thompson5910):

Multiply the corresponding values and add them up: 2*(-1) + 4*5 = -2 + 20 = 18

jimthompson5910 (jim_thompson5910):

So the element in the first row and the first column of the new 2x2 result matrix is 18

OpenStudy (anonymous):

so to get the 2 term in the first row it would be... row 1 times column 2 and add them like you did above. And repeat for row 2?

jimthompson5910 (jim_thompson5910):

yes in a way that's what you're doing to find the element in row 1, column 2 of the answer, you "multiply" row 1 of matrix 1 by column 2 of matrix 2 and add up the products

OpenStudy (anonymous):

18 14 4 5

jimthompson5910 (jim_thompson5910):

2*3 + 4*2 = 6 + 8 = 14 so 14 is in row 1, column 2 of the answer

OpenStudy (anonymous):

Thanks, I was forgetting the adding step.

jimthompson5910 (jim_thompson5910):

you have the correct answer